# Collaborating with AI Agents: A Field Experiment on Teamwork, Productivity, and Performance

### v2

@@ 3.2 Productivity
Participants were incentivized to submit as many ads as possible within the time limit. At the team level, productivity was comparable between human-human and human-AI teams as human-AI teams submitted approximately the same number of ads with half as many workers (Table 5). At the individual level, participants in human-AI teams submitted 73% more ads than their counterparts in human-human teams, with results consistent when demographic controls are added. AI collaboration thus supports individual productivity by enabling participants to generate more output per worker compared to those in human-human teams.
@@
Although our findings suggest that human-AI teams produce a similar amount of output as human-human teams with half as many workers, it is important to note that our study did not include a human-alone condition. Prior studies have consistently shown that AI tools enhance individual productivity compared to humans working alone across various tasks (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Dell’Acqua et al., 2023, Chen and Chan, 2024, Wiles et al., 2023). In light of this evidence, our results suggest that the AI in human-AI teams acts as a near-substitute for an additional human collaborator, enabling a single human to achieve productivity levels comparable to a team with two human workers. Without a human-alone condition, however, we cannot directly quantify AI’s incremental contribution beyond what an individual human might have achieved working alone. We explore this limitation further in the Discussion section.
Submissions
Team-Level Individual-Level
Intercept 5.950∗∗∗ 7.483∗∗∗ 3.203∗∗∗ 3.935∗∗∗
(0.207) (0.832) (0.104) (0.642)
Human-AI -0.406 -0.395 2.341∗∗∗ 2.366∗∗∗
(0.246) (0.246) (0.169) (0.169)
Demographics No Yes No Yes
Observations 1834 1834 2310 2310
  • Notes: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001. Robust standard errors account for heteroskedasticity..
Table 5: Productivity of human-AI and human-human teams.
@@
With that caveat in mind, to test the robustness of our results to the alternative hypothesis that individual-level productivity differences are simply due to the submission of incomplete ads, we examined ad copy completion rates across conditions. As shown in Table 6 and Figure 7, participants in human-AI teams had consistently higher completion rates for ad copy elements compared to participants in human-human teams. Taken together, these findings suggest that human-AI teams are indeed more productive, in both the number of ads submitted and the ad completion rate.
Headline Primary Text Description
Intercept 0.728∗∗∗ 0.737∗∗∗ 0.651∗∗∗
@@
Demographics Yes Yes Yes
Observations 2310 2310 2310
  • Notes: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001. Heteroskedasticity-robust standard errors are reported.
Table 6: Individuals in Human-AI teams submit ads with more copy completed.
Figure 7: Copy completion rates.
3.3 Ad Quality
In addition to the productivity and work process implications of collaborating with AI agents, we also evaluated the performance of human-human and human-AI teams by assessing the quality of their output. Specifically, we examined differences in the quality of ad text and images and conducted field evaluations of ad effectiveness to understand how human-AI collaboration dynamics influence overall effectiveness.
Ad text quality improves, but image quality declines in human-AI teams.
In human evaluations of the quality of the ads, we found that human-AI teams had higher quality ad text but lower quality ad images compared to human-human teams (Table 7, Figure 8). The estimated likelihood of clicking on the ad was indistinguishable between the two groups. Interestingly, in our AI evaluations of ad quality, we found that AI ratings were higher on the text and clicks for ads produced by human-AI teams and the same across the groups for image quality. In a way, it is unsurprising that the AI rated these ads’ text quality, image quality, and estimated click likelihood as equal to or better than those produced by human-human teams because these ads were created in collaboration with OpenAI’s gpt-4o.
However, the human ratings reveal that AI introduces important trade-offs in output quality. Specifically, collaborating with GPT models improves text quality but reduces image quality. This trade-off may arise because GPT models were originally designed for next-word prediction rather than to predict image quality. These results suggest that while GPT models enhance text-based outputs, their contributions to multimodal outputs like ad images may require complementary tools or models designed specifically for image-related tasks. These findings then beg the question of whether and how these trade-offs in text and image quality meaningfully impact actual click through rates, which we examined in detail in the field test of the ads themselves.
Human Evaluations AI Evaluations
Text Image Click Text Image Click
Intercept 4.535∗∗∗ 4.588∗∗∗ 4.730∗∗∗ 4.523∗∗∗ 3.592∗∗∗ 3.433∗∗∗ 5.410∗∗∗ 5.423∗∗∗ 6.438∗∗∗ 6.345∗∗∗ 5.235∗∗∗ 5.246∗∗∗
(0.092) (0.102) (0.080) (0.089) (0.079) (0.088) (0.056) (0.062) (0.054) (0.059) (0.045) (0.051)
Human-AI 0.324∗∗∗ 0.435∗∗∗ -0.134∗∗∗ -0.159∗∗∗ -0.014 0.037 0.122∗∗∗ 0.086∗∗ -0.014 -0.144∗∗∗ 0.068∗∗∗ 0.016
(0.024) (0.053) (0.021) (0.044) (0.021) (0.045) (0.015) (0.031) (0.014) (0.029) (0.013) (0.026)
# Submissions -0.028∗∗∗ -0.012∗∗∗ -0.018∗∗∗ -0.013∗∗∗ -0.010∗∗∗ -0.008∗∗∗
(0.002) (0.001) (0.001) (0.001) (0.001) (0.001)
# AI Images Generated 0.007∗∗∗ 0.029∗∗∗ 0.023∗∗∗ 0.023∗∗∗ 0.028∗∗∗ 0.019∗∗∗
(0.002) (0.002) (0.002) (0.001) (0.001) (0.001)
# Human Messages 0.004∗∗∗ 0.002∗∗ 0.002∗∗∗ 0.003∗∗∗ 0.002∗∗∗ 0.001∗∗∗
(0.000) (0.000) (0.000) (0.000) (0.000) (0.000)
# Human Copy Edits 0.136∗∗∗ 0.042∗∗ 0.082∗∗∗ -0.027 -0.028∗∗ -0.040∗∗∗
(0.017) (0.015) (0.015) (0.011) (0.010) (0.010)
# Human Copy Edits $\times$ (H-AI) -0.033 -0.004 -0.022 -0.020 0.028 -0.048∗∗
(0.027) (0.027) (0.026) (0.018) (0.017) (0.016)
# Human Image Edits -2.173 -1.019 -0.560 -2.165∗∗∗ -2.390∗∗∗ -1.853∗∗∗
(0.897) (0.707) (0.726) (0.542) (0.459) (0.442)
# Human Image Edits $\times$ (H-AI) 2.057 1.974 1.665 1.057 2.789∗∗∗ 1.555∗∗
(1.009) (0.829) (0.832) (0.591) (0.514) (0.487)
Demographics Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
Observations 45803 45803 45803 45803 45803 45803 11138 11138 11138 11138 11138 11138
  • Notes: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001. Standard errors are clustered at the ad level to account for within-ad correlation. “H-AI” refers to human-AI teams. “Human AI Images Generated” refers to the AI images generated by humans. Copy and image edits are per thousand.
Table 7: Human and AI ratings of ads.
Figure 8: AI and human ratings of ads.
We also examined how work process variables, like the number of messages exchanged, the number of ads submitted, the number of AI images generated, and the number of copy and image edits, affect ad quality in post hoc analyses. We found that more messages exchanged, fewer ads submitted, and more AI images generated were associated with higher human- and AI-evaluated ad quality. Interestingly, for text copy and image edits, we observed nuanced differences in ad quality across team types and human and AI evaluations. Our analyses showed that humans rate ads more highly when humans copy edit text more in human-human teams. In contrast, AI rate ads more highly when humans copy edit text less in human-human teams. Human text copy edits made in human-AI teams mostly revealed no difference in preference in either human or AI evaluations of ad quality (except for a slight negative preference for AI-evaluated click quality). Both humans and the AI rated ads higher when image edits were made in human-AI teams and rated ads lower when image edits were made in human-human teams. This finding is consistent with previous our findings that humans are better at selecting images—as human edits to images are more highly valued, by both humans and AI, in teams in which AI is contributing to the image selection and generation process, humans seem to be needed to correct, adjust or improve image selection and generation when AI is involved. These findings highlight that team composition (human-human vs human-AI teams) and work processes (messages exchanged, ads submitted, AI images generated, and copy and image edits) all meaningfully shape ad quality.
3.4 Ad Effectiveness
To evaluate the real-world performance of ads created by human-human and human-AI teams, we conducted a field experiment on the social media platform X. Our ad campaigns generated 4,932,373 impressions and 7,546 clicks over 20 days. This section examines how collaboration type influences key advertising metrics including the cost-per-click (CPC) and click-through rate (CTR), reported in Table 8, and the view-through rate (VTR) and view-through duration (VTD) shown in Table 9. The field study extends lab findings and tests how the distinct productivity and quality profiles of human-AI and human-human teams translate to advertising outcomes in a real world setting. Broadly, we found that ads with higher text quality, created more often by human-AI teams, and higher image quality, created more often by human-human teams, performed significantly better on cost per click and click through rate metrics. As human-AI teams produced higher quality text and lower quality images, their effects counteracted one another and human-AI ads performed similarly to those created by human-human teams overall.
@@ 3.4 Ad Effectiveness
We examined CPC and CTR using regression models with campaign random effects to account for unobserved heterogeneity across the 400 campaigns, with results in Table 8. For CPC, measured in dollars, our analysis in Column 1 reveals no significant effect of collaboration type. [^6] Higher spend consistently lowers costs and suggest divergent targeting and optimization, likely because we had kept on auto-bidding in our campaigns. Thus, we consider the effects present in this section as conditional on ad algorithms, as per recommendation 5 from Braun et al. (2024). With human-evaluated quality scores (i.e., Image, Text, Click) as covariates, ads with stronger image quality significantly reduced CPC by \$0.268 or 2.3% (Column 2), and this pattern held when all predictors were included in the analysis (Column 3). For CTR, expressed as a percentage, the analysis with collaboration indicated no direct human-AI effect (Column 4). With quality scores, we found that text quality significantly reduced CTR by 0.008% (Column 5), which remained consistent when all factors were included (Column 6).
CPC (\$) CTR (%)
Intercept 10.234∗∗∗ 11.636∗∗∗ 11.560∗∗∗ 0.000 -0.027 -0.026
(0.323) (0.686) (0.698) (0.006) (0.013) (0.013)
Human-AI 0.115 0.128 -0.000 -0.001
(0.208) (0.210) (0.004) (0.004)
Image -0.268 -0.264 0.002 0.002
(0.131) (0.131) (0.002) (0.002)
Text -0.154 -0.163 0.008∗∗∗ 0.008∗∗∗
(0.114) (0.115) (0.002) (0.002)
Click 0.212 0.215 -0.002 -0.002
(0.147) (0.147) (0.003) (0.003)
Spend -0.135∗∗∗ -0.141∗∗∗ -0.140∗∗∗ 0.006∗∗∗ 0.006∗∗∗ 0.006∗∗∗
(0.012) (0.013) (0.013) (0.000) (0.000) (0.000)
Campaign RE Yes Yes Yes Yes Yes Yes
Observations 1859 1859 1859 2000 2000 2000
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Campaign RE represents campaign random effects. Ads with zero clicks were removed for regressions for columns 1-3.
Table 8: Effects on cost-per-click (CPC) and click-through rates (CTR) from the field study.
@@
We assessed view metrics—view-through rate (VTR) and view-through duration (VTD; in log-seconds)—using similar regression models, as shown in Table 9. Our analysis of VTD followed a similar pattern as observed in the click metrics—while there was no direct human-AI effect (Column 4), text quality was significantly associated with longer viewing duration. On average, one point higher text quality ratings were associated with one second longer view through duration (Columns 5-6). Our analysis of VTR showed no significant effects for collaboration type or quality (Columns 1-3). These findings demonstrate that ads with higher text quality, which are more often produced by human-AI teams, are viewed for longer periods of time.
VTR VTD (log-sec)
Intercept 0.000 0.000 0.000 0.509∗∗∗ 0.000 0.000
(0.006) (0.011) (0.012) (0.052) (0.116) (0.118)
Human-AI 0.001 0.000 0.021 0.011
(0.004) (0.004) (0.037) (0.037)
Click -0.000 -0.000 -0.003 -0.003
(0.003) (0.003) (0.026) (0.026)
Image -0.003 -0.003 -0.001 -0.000
(0.002) (0.002) (0.023) (0.023)
Text 0.002 0.002 0.038 0.037
(0.002) (0.002) (0.019) (0.019)
Spend 0.001∗∗∗ 0.001∗∗∗ 0.001∗∗∗ -0.001 0.005 0.005∗∗
(0.000) (0.000) (0.000) (0.002) (0.002) (0.002)
Campaign RE Yes Yes Yes Yes Yes Yes
Observations 2000 2000 2000 2000 2000 2000
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Campaign RE represents campaign random effects.
Table 9: Effects on view-through rate (VTR) and view-through duration (VTD) from the field study.

### v3

@@ 4.1 Productivity, Quality and Output Diversity
Participants were incentivized to submit as many ads as possible within the time limit. Participants in human-AI teams submitted 50% more ads per worker, per unit time than their counterparts in human-human teams, with results consistent when demographic controls are added (Table 2). AI collaboration thus supports individual productivity by enabling participants to generate more output per worker compared to those in human-human teams.
Productivity Quality Output
Text Image Click Diversity
Intercept 4.601∗∗∗ 5.011∗∗∗ 4.666∗∗∗ 3.698∗∗∗ 0.303∗∗∗
(0.346) (0.038) (0.034) (0.035) (0.010)
Human-AI 1.954∗∗∗ 0.327∗∗∗ -0.133∗∗∗ -0.013 -0.094∗∗∗
(0.197) (0.024) (0.021) (0.021) (0.006)
Demographics Yes Yes Yes Yes Yes
Observations 2234 45803 45803 45803 2034
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Productivity is the number of submissions by participants. Text, Image, Click are human ratings (7-point scale) with standard errors clustered at the ad level. Output Diversity is cosine distance to treatment centroid (higher is more diverse). Robust standard errors account for heteroskedasticity. Full ratings model with covariates appears in Table G.
Table 2: Collaborating with AI agents improves productivity and text quality but lowers image quality (i.e., a jagged frontier), and decreases output diversity.
@@
It is important to note that our study did not include a human-alone condition. Prior studies have consistently shown that AI tools enhance individual productivity compared to humans working alone across various tasks (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Dell’Acqua et al., 2023, Chen and Chan, 2024, Wiles et al., 2023). In light of this evidence, our results suggest that the AI in human-AI teams acts as a partial substitute for an additional human collaborator. Without a human-alone condition, however, we cannot directly quantify AI’s incremental contribution beyond what an individual human might have achieved working alone. We explore this limitation further in the Discussion section.
@@
With that caveat in mind, to test the robustness of our results to the alternative hypothesis that individual-level productivity differences are simply due to the submission of incomplete ads, we examined ad copy completion rates across conditions. As shown in Table E, participants in human-AI teams had consistently higher completion rates for ad copy elements compared to participants in human-human teams. Taken together, these findings suggest that human-AI teams are indeed more productive, in both the number of ads submitted and the ad completion rate.
Collaborating with AI agents improves text quality but lowers image quality.
In human evaluations of the quality of the ads, we found that human-AI teams produced higher quality ad text but lower quality ad images compared to human-human teams (Table 2). The estimated likelihood of clicking on the ad was indistinguishable between the two groups. The human ratings reveal that AI introduces important trade-offs in output quality. Specifically, collaborating with LLM-based agents improves text quality but reduces image quality. This pattern aligns with the jagged frontier of AI capabilities (Dell’Acqua et al., 2023), where AI excels at some tasks (text generation) while underperforming at others (image quality prediction). These results were supported by AI evaluations of ad quality (see Table H) and may arise because LLMs are trained for next-word prediction rather than image quality prediction. [^5] Interestingly, in our AI evaluations of ad quality, we found that AI ratings were higher on the text and clicks for ads produced by human-AI teams and the same across the groups for image quality. In a way, it is unsurprising that the AI rated these ads’ text quality, image quality, and estimated click likelihood as equal to or better than those produced by human-human teams because these ads were created in collaboration with OpenAI’s gpt-4o. These results suggest that while LLMs enhance text-based outputs, their contributions to multimodal outputs like ad images may require complementary tools or fine-tuning for image-related tasks.
Collaborating with AI agents reduces output diversity.
Beyond productivity and quality, we examined whether collaboration with AI agents affects output diversity. Using cosine distance to treatment centroids as a measure of output diversity (Table 2), human-AI teams produce outputs that are closer to their condition centroid than human-human teams ($\beta=-0.094$, $p<0.001$), indicating a homogenization of teams outputs (Chen and Chan, 2024, Padmakumar and He, 2024). This diversity collapse suggests that while AI-generated content may be higher quality on average on certain dimensions, it lacks the variation that human creativity provides.
4.2 Ad Performance in the Field
Our findings on productivity and quality raise an important question about whether and how these trade-offs meaningfully impact actual performance. To evaluate the real-world performance of ads created by human-human and human-AI teams, we conducted a field experiment on the social media platform X. Our ad campaigns generated 4,932,373 impressions and 7,546 clicks over 20 days. This section examines how human-AI collaboration influences key advertising metrics, including cost-per-click (CPC), click-through rate (CTR), and view-through duration (VTD), reported in Table 3, and view-through rate (VTR) shown in Table I. The field study extends the lab experiment findings and tests how the distinct productivity and quality profiles of human-AI and human-human teams translate to advertising outcomes in a real world setting. Broadly, we found that image quality reduces CPC while text quality increases CTR and VTD. Since human-AI teams produce higher text quality but lower image quality than human-human teams, these effects are offset, and human-AI ads perform similarly to human-human ads overall.
CPC (\$) CTR (%) VTD (log-sec)
Intercept 11.628∗∗∗ -0.028 0.000
@@ 4.2 Ad Performance in the Field
Image -0.290 0.002 -0.002
(0.132) (0.003) (0.023)
Text -0.149 0.007∗∗∗ 0.037
(0.116) (0.002) (0.019)
Click 0.209 -0.002 -0.008
(0.148) (0.003) (0.027)
Spend Yes Yes Yes
Campaign RE Yes Yes Yes
Observations 1822 1962 4997
  • Notes: p<<0.1, p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Spend is controlled to account for auto-bidding. Campaign RE represents campaign random effects. CPC excludes ads with zero clicks. Robust standard errors account for heteroskedasticity.
Table 3: In the field experiment, higher image quality reduces cost-per-click while higher text quality increases click-through rate and view-through duration.
@@
We examined CPC and CTR using regression models with campaign random effects to account for unobserved heterogeneity across the 400 campaigns, with results in Table 3. For CPC, measured in dollars, our analysis reveals no significant effect of collaboration type. [^6] Higher spend consistently lowers costs and suggests divergent targeting and optimization, likely because we had kept on auto-bidding in our campaigns. Thus, we consider the effects present in this section as conditional on ad algorithms, as per recommendation 5 from Braun et al. (2024). Ads with stronger image quality significantly reduced CPC by \$0.29 or 2.5%. For CTR, expressed as a percentage, there was no direct human-AI effect. Text quality increased CTR by a small but statistically significant 0.007%. Thus, quality trade-offs between collaboration types translate to distinct click performance advantages.
@@
We assessed view-through rate (VTR) and view-through duration (VTD; in log-seconds), using similar regression models, with VTD in Table 3 and VTR in Table I. Our analysis of VTD followed a similar pattern as observed in the click metrics. While there was no direct human-AI effect, text quality was associated with longer viewing duration ($p$ = 0.051). On average, one point higher text quality ratings were associated with approximately 4% longer view-through duration. Our analysis of VTR showed no significant effects for collaboration type or quality. These findings demonstrate that ads with higher text quality, which are more often produced by human-AI teams, are viewed for longer periods of time.
4.3 Teamwork and Collaboration Dynamics
Having established the outcomes of human-AI collaboration, we now examine differences in the collaboration process itself. Teams collaborate through two primary modes: communication and task execution. We analyze both to understand the shifts in teamwork dynamics underlying the productivity, quality and performance outcomes.
Collaborating with AI agents increases overall communication.
To test how human-AI collaboration affects communication, we measured the number of messages sent by participants in both human-human and human-AI groups. Participants worked collaboratively with and sent messages to their partners in real-time, for both human-human and human-AI groups. In the human-human group, participants messaged their human partners; in the human-AI group, participants messaged the AI. The Pairit platform logs a timestamped record of each message sent by a participant or AI. Participants in human-AI teams sent 62% more messages than those in human-human teams, with consistent results when demographic covariates were included (Table 4). These results indicate that collaborating with an AI partner encourages more frequent communication overall.
Communication Task Execution
Count Task-Oriented Interpersonal Copy Edits Delegation
Intercept 20.296∗∗∗ 0.362∗∗∗ 0.580∗∗∗ 1676.458∗∗∗ 0.478∗∗∗
(1.450) (0.014) (0.015) (57.646) (0.018)
Human-AI 12.580∗∗∗ 0.090∗∗∗ -0.107∗∗∗ -1032.406∗∗∗ 0.083∗∗∗
(0.843) (0.008) (0.008) (33.588) (0.010)
Demographics Yes Yes Yes Yes Yes
Observations 2234 2234 2234 2234 2233
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Count is the number of messages sent by a participant. Task-oriented is the fraction of messages labeled as content or process. Interpersonal is the fraction of messages labeled as emotional or social. Copy edits are counts of text changes (i.e., text input events). Delegation is fraction of text work performed by partner. Robust standard errors account for heteroskedasticity.
Table 4: Collaborating with AI agents increases communication, with more task-oriented and less interpersonal messages. It also decreases direct edits to the ad copy while increasing the delegation to the AI agents.
Collaborating with AI agents shifts teams from interpersonal toward task-oriented communication.
In addition to the number of messages sent across conditions, we investigated whether the content of messages varied across conditions. We used gpt-4o-mini to label each message into categories, which we grouped into task-oriented and interpersonal communication (see Section 3.7). We found that human-AI teams sent 25% more task-oriented messages, while human-human teams sent 18% more interpersonal messages (Table 4). This shift indicates that collaboration with AI emphasizes task-oriented over interpersonal interaction, possibly because participants can focus more on the task without needing to navigate the social or emotional aspects typical of human collaboration.
Collaborating with AI agents shifts work from direct editing to delegation.
Participants in human-AI teams made 62% fewer direct copy edits compared to human-human teams (Table 4). However, participants in the human-AI condition still engaged in direct editing, indicating that while AI collaboration reduced the frequency of manual edits, it did not eliminate them entirely. Human-AI teams also delegated 58% of work to AI partners compared to 50% in human-human teams ($\beta=0.084$, $p<0.001$), a 17% increase. Unlike previous studies, such as Chen and Chan (2024), which artificially constrain participants to specific modalities when interacting with AI, our results show that in more unconstrained collaborative settings, where humans and agents can freely allocate effort to different tasks, participants utilized both direct involvement with the work product (ad editing) and interaction through messaging.
Taken together, these patterns in communication and task execution support a delegation workflow: humans direct AI partners rather than collaborating as peers. Human-AI teams communicate more frequently and with greater emphasis on task-oriented communication, while making fewer direct edits and delegating more work to their AI partners.
4.4 Mechanisms: Teamwork and Collaboration Dynamics Moderate Performance Outcomes
To address why collaborating with AI agents produces these outcomes, we conducted mediation analyses explained in Section 3.10. We found that participants who correctly recognized their AI partner engaged in more task-oriented and less interpersonal communication and delegate more work to their AI partner. We also found that directive communication was positively associated with more delegation and that both communication style and delegation predict output quality.
First, participants in the human-AI treatment group were highly accurate in identifying their partner as AI (94.5%), while human-human group participants were relatively less accurate (57.6%). When we tested whether this recognition shaped collaboration behavior (Equation 3) we found that correct identification predicts communication volume, style, and task execution (Table 5). Participants who correctly identified their partner as an AI agent sent more messages ($\beta_{3}=11.453$, $p<0.001$), with a higher proportion of task-oriented messages ($\beta_{3}=0.061$, $p<0.001$) and a lower proportion of interpersonal messages ($\beta_{3}=-0.059$, $p<0.01$). They also delegated more ($\beta_{3}=0.056$, $p<0.05$) and, despite making far fewer direct copy edits overall in human-AI teams ($\beta_{1}=-1155$, $p<0.001$), made slightly more edits than participants who did not correctly identify their partner ($\beta_{3}=150.6$, $p<0.05$), suggesting that recognition may prompt greater oversight of AI-generated content. Together, these findings indicate that awareness of AI capabilities prompts both a shift toward task-oriented, directive communication, more delegation, and a calibrated approach to task execution with greater oversight.
Communication Task Execution
Count Task-Oriented Interpersonal Copy Edits Delegation
Intercept 17.069∗∗∗ 0.371∗∗∗ 0.584∗∗∗ 1632.538∗∗∗ 0.483∗∗∗
(1.552) (0.017) (0.017) (69.997) (0.020)
Human-AI $\times$ 11.453∗∗∗ 0.061∗∗∗ -0.059∗∗ 150.632 0.056
   Recognition (1.676) (0.018) (0.019) (66.080) (0.023)
Human-AI 3.280 0.044∗∗ -0.066∗∗∗ -1155.053∗∗∗ 0.043
(1.317) (0.016) (0.016) (48.054) (0.020)
Recognition 3.904∗∗∗ -0.011 0.007 53.490 -0.012
(0.997) (0.013) (0.013) (61.577) (0.011)
Demographics Yes Yes Yes Yes Yes
Observations 2234 2228 2228 2234 2233
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Recognition indicates whether the participant correctly identified their partner type. Count is the number of messages sent by a participant. Task-oriented is the fraction of messages labeled as content or process. Interpersonal is the fraction of messages labeled as emotional or social. Copy edits are counts of text changes (i.e., text input events). Delegation is fraction of text work performed by partner. Robust standard errors account for heteroskedasticity.
Table 5: Recognition of AI partner moderates communication and task execution.
Second, communication style and delegation predict productivity, ad quality and output diversity, but these effects vary across human-AI and human-human teams (Table 6, Panel A). In human-human teams, task-oriented communication is negatively associated with productivity ($\beta=-4.19$, $p<0.001$), text quality ($\beta=-0.416$, $p<0.05$), and image quality ($\beta=-0.267$, $p<0.05$). However, in human-AI teams, the reverse is true: task-oriented communication improves productivity ($\beta=4.69$, $p<0.001$) and ad quality when working with AI ($\beta=0.746$, $p<0.001$ for text; $\beta=0.506$, $p<0.01$ for image; $\beta=0.584$, $p<0.001$ for click). In contrast, while interpersonal communication is positively associated with productivity in human-human teams ($\beta=4.25$, $p<0.001$), it is negatively associated with productivity ($\beta=-4.80$, $p<0.001$) and ad quality in human-AI teams ($\beta=-0.685$, $p<0.001$ for text; $\beta=-0.581$, $p<0.001$ for image; $\beta=-0.599$, $p<0.001$ for click). These patterns suggest that interpersonal communication facilitates productivity in human-human teams, perhaps through rapport-building and coordination, whereas task-oriented, directive communication is more effective in human-AI teams.
Productivity Quality Output
Text Image Click Diversity
Panel A: Communication
Task-Oriented -4.187∗∗∗ -0.416 -0.267 -0.286 0.011
(0.847) (0.166) (0.126) (0.124) (0.027)
Task-Oriented $\times$ H-AI 4.693∗∗∗ 0.746∗∗∗ 0.506∗∗ 0.584∗∗∗ -0.035
(1.242) (0.206) (0.173) (0.166) (0.035)
Interpersonal 4.247∗∗∗ 0.269 0.211 0.194 0.014
(0.754) (0.156) (0.118) (0.119) (0.026)
Interpersonal $\times$ H-AI -4.799∗∗∗ -0.685∗∗∗ -0.581∗∗∗ -0.599∗∗∗ 0.030
(1.141) (0.195) (0.168) (0.161) (0.034)
Demographics Yes Yes Yes Yes Yes
Observations 2041 2041 2041 2041 2030
Panel B: Task Execution
Delegation -1.382 0.302∗∗∗ 0.132 0.178∗∗ -0.083∗∗∗
(0.586) (0.071) (0.069) (0.065) (0.012)
Demographics Yes Yes Yes Yes Yes
Observations 1222 1222 1222 1222 1222
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Panel A: both treatments; base coefficients are H-H. Panel B: H-AI only. Task-oriented is the fraction of messages labeled as content or process. Interpersonal is the fraction of messages labeled as emotional or social. Delegation is fraction of text work performed by partner. Productivity is the number of submissions by participants. Text, Image, Click are human ratings (7-point scale) with standard errors clustered at the ad level. Output Diversity is cosine distance to treatment centroid (higher is more diverse). Robust standard errors account for heteroskedasticity.
Table 6: Communication style and task execution predicts performance differently in human-AI vs. human-human teams.
In human-AI teams, delegation improves text ratings ($\beta=0.302$, $p<0.001$) and click ratings ($\beta=0.178$, $p<0.01$) but has no effect on image ratings ($\beta=0.132$, n.s.; Table 6, Panel B), consistent with the jagged frontier of AI capabilities (Dell’Acqua et al., 2023). Click ratings likely track text quality because clicking is partially driven by copy appeal, and delegation improves text without worsening image quality. Delegation also reduces output diversity ($\beta=-0.083$, $p<0.001$), explaining the diversity collapse observed in human-AI teams. Notably, delegation is associated with lower productivity ($\beta=-1.38$, $p<0.05$), perhaps because excessive delegators are less engaged overall. In summary, recognition shifts communication style and increases delegation, which in turn affects ad quality and diversity.
Illustrative examples.
To ground these quantitative findings, we extracted representative message exchanges. In H-AI teams, communication is directive: “Can you generate an image with a professional looking person at a desk?” followed by the AI responding with a generated image. Coordination appears as explicit task allocation: “Let’s divide this up. I’ll work on the headline and you handle the image.” In contrast, H-H teams show rapport-building: “Sorry for the delay! Had to grab coffee” … “No worries, take your time.” H-H teams also engage in collaborative negotiation: “What do you think about using blue instead?” … “Hmm, I was thinking green might work better because…” … “Good point, let’s try both and see.” These exchanges illustrate the directive versus collaborative dynamics that distinguish human-AI from human-human teams.